OpenStudy (anonymous):

Solve the quadratic expression by completing the square. x2 + 10x – 17 = 7 I did the problem but when I plug my answers into the original equation it doesn't work. this is what i got. x^2+10x-24=0 (x+2)(x-12)=0 x=-2,12

OpenStudy (anonymous):

You need to change the signs...(x + 2)(x -12) = x^2 - 10x -24!

OpenStudy (anonymous):

OHH I see, and wait do I have to use some other method to solve this, since it says to solve by completing the square. I have to show my work so I have to make sure I do the problem the way it says or I will get it wrong.

OpenStudy (anonymous):

Good point...you do need to use a different method. See this link: http://www.purplemath.com/modules/sqrquad.htm

hero (hero):

Yes, he should learn how to complete the square because x^2 - 10x - 24 is not even the right polynomial to use.

OpenStudy (anonymous):

Can I please get some help on how to do this the right way? I'm confused

OpenStudy (anonymous):

I think I probably know how to complete the square, I just need to re-jog my memory a bit

hero (hero):

\[x^2+10x-24=0 \\x^2 + 10x = 24 \\x^2 + 10x + 25 = 24 + 25 \\(x + 5)^2 = 49 \\x + 5 = \pm 7 \\x = \pm7 - 5 \\x = 2 \\x = -12\]

hero (hero):

The trick is to remember to get whatever you're working with to the form \(ax^2 + bx = c\) Then afterwards, add \(\left(\frac{b}{2}\right)^2\) to both sides.

OpenStudy (anonymous):

Thanks so much, I get it:)